Q 7.127:

A single force acts on a 4.3 kg4.3 kg particle-like object in such a way that the position of the object as a function of time is given by x = 4.2t + 3.1t2 + 2.2t3x = 4.2t + 3.1t2 + 2.2t3 , with xx in meters and tt in seconds. Find the work done on the object by the force from t = 1.4 st = 1.4 s to  t = 3.5 st = 3.5 s .

The answer is 23067
Answer #1

given

mass m = 4.3 kg

x = 4.2 t + 3.1 t2 + 2.2 t3  

t = 1.4 sec to 3.5 sec

v = dx/dt = d ( 4.2 t + 3.1 t2 + 2.2 t3 ) / dt

v = 4.2 + 6.2 t + 6.6 t2

v1.4 = 4.2 + 6.2 X 1.4 + 6.6 X 1.42

v1.4 = 25.81 m/sec

sameway

v = dx/dt = d ( 4.2 t + 3.1 t2 + 2.2 t3 ) / dt

v = 4.2 + 6.2 t + 6.6 t2

v3.5 = 4.2 + 6.2 X 3.5 + 6.6 X 3.52

v3.5 = 106.75 m/sec

W = 1/2 m ( V3.52 - V1.42 )

W = 0.5 X 4.3 X ( 106.752 - 25.812 )

W = 2.15 X ( 10729.4064 )

W = 23068.22 J

work done on the object by the force from t = 1.4 s to  t = 3.5 s is W = 23068.22 J